OpenStudy (anonymous):

Scores on a dental anxiety scale range from 0 (no anxiety) to 20 (extreme anxiety). The scores are normally distributed with a mean of 11 and a standard deviation of 4. Find the z-score for the following raw score: 12

OpenStudy (anonymous):

A z-score (aka, a standard score) indicates how many standard deviations an element is from the mean. A z-score can be calculated from the following formula. z = (X - μ) / σ where z is the z-score, X is the value of the element, μ is the population mean, and σ is the standard deviation. Taken from http://stattrek.com/statistics/dictionary.aspx?definition=z_score.

Directrix (directrix):

z-score = (raw score - mean)/ standard dev z-score = (12 - 11) / 4 = 1/4 or .25

OpenStudy (anonymous):

thank you!